Circular No. 5738

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 W. Wamsteker, P. M. Rodriguez and R. Gonzalez, ESA IUE Obser-

vatory, for the ESA IUE ToO team for Supernovae; G. Sonneborn, NASA Goddard Space Flight Center; and R. Kirshner, Center for Astrophysics, report: “Low-resolution IUE spectra (0.5 nm) covering 115-320 nm show a strong ultraviolet continuum over the whole range, with a sharp dropoff shortward of 127.5 nm. A weak 220-nm feature suggests total extinction of only E(B − V) = 0.07. The short-wavelength fluxes decreased by ∼ 30-40 percent between Mar. 30.2 and 30.8 UT; no decrease was evident in the fluxes longward of 240 nm. Many interstellar absorption lines are seen in the spectrum attributable to Si II + O I 130.0, C II 133.5, Si IV 140.0, C IV 155.0, Al II 167.1, Fe II + Mn II 260.0, Mg II 280.0 and Mg I 285.2 nm. These features are considerably weaker than seen in the similar spectra of SN 1980K with the same instrument. The flux F (in 5-nm bands) in units of 10^{−14} erg s^{−1} cm^{−2} A^{−1} on Mar. 30.2 was F (127.5) = 126, F (173.5) = 115, F (245.0) = 75, F (290.0) = 68. High-resolution spectra (0.02 nm) show a complex structure of the numerous interstellar lines. The short-wavelength spectrum also shows the emission features N V 123.88 and 124.28 nm redward of rest wavelength in the frame of M81. An LSR velocity of –100 km/s and M81 velocity of –40 km/s were assumed. N V 123.88 nm has an FWHM < 50 km/s; 124.28 nm has a red wing extending > 200 km/s from line center. Low-ionization interstellar absorption lines such as Mg I 285.2 nm show a multi-component structure at –117, –63, +12, +129 and +181 km/s (heliocentric); various weaker components may also be present.”

COMET SHOEMAKER-LEVY (1993e) The following prediscovery observations have been reported:

1993 UT α_{2000} δ_{2000} m_{1} Observer
Mar. 17.75868 12^{h}29^{m}50.^{s}80 −4^{o}22′13′′.2 14.0 Otomo
17.77153 12 29 50.21 −4 22 12.4  _{”}
19.39514 12 29 03.41 −4 17 43.0 12  Helin
19.42014 12 29 02.78 −4 17 36.3  _{”}
21.46788 12 28 03.61 −4 11 50.7  _{”}

S. Otomo (Otomo Observatory). 0.25-m f/3.4 reflector. E. Helin, K. Lawrence and C. Brewer (Palomar). 0.46-m Schmidt. Mea-

surer Lawrence. The Mar. 19 images were actually noted at the time. Corrigendum. On IAUC 5725, the magnitude 16.7 should be an m_{2}.

1993 April 1 Brian G. Marsden